# Localizing Events in Videos with Multimodal Queries

### v1

@@ 4.2 Results & Analysis
Captioning
@@
Visual Query Enc.
Moment-DETR (2021) 12.55 5.69 13.38 6.59 14.36 6.01 14.88 6.53
QD-DETR (2023) 15.91 9.12 14.88 8.62 13.90 8.49 14.62 8.36
QD-DETR (2023) 15.65 10.03 12.60 6.79 12.34 6.72 12.34 7.44
EaTR (2023) 19.86 13.00 19.91 12.99 21.15 13.45 21.48 13.38
CG-DETR (2023) 22.90 13.00 24.93 13.58 23.24 13.12 24.74 14.23
TR-DETR (2024) 17.92 11.19 17.36 11.10 15.14 9.86 15.60 9.53
UMT (2022) 5.43 2.85 4.77 2.09 5.22 2.35 4.57 2.42
UniVTG (2023) 21.93 13.00 23.89 13.64 22.78 13.19 22.52 12.79
UVCOM (2023) 17.08 9.77 16.78 10.97 17.36 11.68 17.10 11.23
{NiceTabular}llcccccccc[colortbl-like]
[…]
9-11
1
[…]
scribble
cartoon
cinematic
realistic
mAP@0.5 Avg. mAP@0.5 Avg. mAP@0.5 Avg. mAP@0.5 Avg.
Captioning
Summarization
Visual Query Enc.
Moment-DETR (2021) 14.95 6.67 16.51 7.21 17.00 7.39 17.41 7.66
QD-DETR (2023) 19.48 10.11 19.57 10.18 18.07 9.54 18.88 9.94
QD-DETR (2023) 18.22 9.74 14.31 7.30 15.18 7.45 14.71 7.66
EaTR (2023) 25.27 13.98 25.95 14.21 26.83 14.70 26.65 14.49
CG-DETR (2023) 30.24 15.57 30.78 15.70 30.07 15.48 30.98 15.83
TR-DETR (2024) 21.09 11.67 20.87 11.71 19.62 11.02 19.72 10.76
UMT (2022) 5.57 2.81 4.66 1.96 5.60 2.46 4.59 2.23
UniVTG (2023) 24.30 13.02 20.80 11.56 19.85 10.99 19.42 10.95
UVCOM (2023) 20.13 11.15 20.19 11.96 20.67 12.37 20.73 12.03
Best adaptation methods We find that Cap can achieve the best performance and is more robust to different prompts compared to other adaptation methods by an average margin of 3.6% on all styles. We observe that both utilizing MLLMs for captioning reference images, Sum suffers more than Cap adaptation regarding performance and is more sensitive to prompts for all reference styles, which can be observed from the higher standard deviation, showing asking MLLMs to caption and summarize the refinement texts is less controllable. To conclude, captioning images is still a golden method since MLLMs and LLMs are powerful enough to generate faithful captions.
Comparing models Models that perform well in one adaptation method tend to perform well in others. For example, UVCOM and TR-DETR consistently show high performance across Cap, Sum, and VisEnc methods. We observe that more recent models keep their outperforming performance on our ICQ. Latest models, including UVCOM, TR-DETR, and CG-DETR, tend to perform better across different adaptation methods and reference image styles. In contrast, older models like Moment-DETR consistently show lower performance. LLM-based models, SeViLA, cannot compete with other specialized models; this aligns with their subpar performance on natural language query-based benchmarks. In the next section, we find that model performance on ICQ highly correlates with that on natural language query-based benchmark QVHighlight. This shows that (1) our multimodal queries share semantics with the original benchmark; (2) the adaptation methods and models could understand semantics from multimodal queries.
Comparing styles We find all adaptation methods perform consistently across different styles and therefore suggest that they could understand the multimodal semantics queries well, particularly for styles including cartoon, cinematic, and realistic; the model performance is close to each other. For scribble, all models show marginally worse performance, and even both Cap and Sum methods have more significant standard deviation, which reflects that it is heavily influenced by the prompts. This can be explained by the fact that scribble images are more minimal and abstract in semantics and more challenging to interpret. Surprisingly, in spite of being more abstract and simpler, the model performance on scribble reference images is close to other reference image styles. This demonstrates the potential of using scribble as multimodal queries in real-world video event localization applications like video search.
Comparing refinement text types In addition, we calculate the model performance on different subsets of refinement texts shown in Fig. 5. We conclude even though models have close performance across reference image styles, and they show varied performance on different refinement text types across styles. For scribble style, models perform in general for “relation” better than on other styles. For cartoon style, models demonstrate a more balanced performance across all types. The performance is notably higher for “environment” and “attribute” in cinematic style. Finally, for realistic style, the models yield better performance in “object” and “environment”.
4.3 Ablation studies
Multimodal Query-based vs. Natural Language Query-based Performance We compare model performance on the multimodal query-based ICQ-Highlight and the original natural language query-based QVHighlight using Spearman’s rank correlation coefficient [53] on R1@0.5. For scribble, Spearman’s rank correlation coefficients are 0.89(Cap) and 0.93(Sum). The cartoon style yields coefficients of 0.98(Cap) and 0.94(Sum). The cinematic style shows coefficients of 0.93 for both Cap and Sum. Lastly, realistic has coefficients of 0.96(Cap) and 0.95(Sum). The high correlation scores indicate a strong positive correlation across benchmarks, suggesting queries of both benchmarks share the common semantics and yield the reliability of our benchmark.
Model Performance With vs. Without Refinement Texts To assess the impact of refinement texts on video event localization using multimodal queries, we have evaluated model performance using only reference images as queries, omitting refinement texts. We employ the Cap adaptation without a modifier for integrating refinement texts. As shown in Tab. 3, we present the model performance and their relative performance drop in percentage compared to those with refinement texts. Models have different scales of performance drop, which indicates that refinement texts help refine the semantics of reference images and localize the events. Additionally, we observe that for scribble images, the performance drop is less pronounced compared to other reference image styles in that these images are inherently minimalistic and less reliant on detailed semantics.
@@ 5 Conclusion
Limitations and Future Work As the first multimodal query-based video event localization benchmark, ICQ still has several limitations: (1) Selection of LLM-based models is limited as of the time of the work due to the lack of more open source models; (2) Our current benchmark utilizes generated multimodal queries and, as a result, can suffer from generation artifacts. Additionally, although our benchmark serves as a practical testbed, fine-tuning models with unlabeled videos [43, 59] for this new setting remains an open question, particularly because of the lack of training data.
Societal Impacts We believe that using multimodal semantic queries for video event localization brings prospects in real-world applications, such as providing service for illiterate, pre-literate, or non-speakers in cross-lingual situations, as it allows them to interact with videos through simple scribble images as a more accessible and convenient approach. However, reference images could contain intentionally harmful content, which may pose new threats to AI safety and privacy.
In this work, we introduce a new benchmark, ICQ, marking an initial step towards using multimodal semantic queries for video event localization. We have found that MLLM/LLM-enhanced enhanced adaptation methods can accommodate conventional models to multimodal queries, serving as a simple yet effective baseline for this novel setting. Our findings confirm that using multimodal queries for video event localization is practical and feasible. Nonetheless, the field remains open to innovative model architectures and training paradigms for multimodal queries. We believe our work paves the way for real-world applications that leverage multimodal queries to interact with video content.
Appendix A Appendix
@@
Additional information about the dataset ICQ-Highlight and licenses for the datasets and models we have used.
@@
A.1 Dataset: ICQ-Highlight
A.1.1 License
@@
A.1.2 Construction Pipeline
We base our model on the original annotation from QVHighlights [29]. The whole pipeline as shown in Fig. 6 consists of (1) annotation: We further conduct a quality check on the annotations in the original dataset and filter out a few samples (details can be found in Sec. A.1.4). In order to generate more relevant reference images, we manually augment the original captions by adding new visual details based on three frames extracted from the raw videos. To introduce refinement texts, we purposely alter certain details of the captions to generate a new one. All annotations are carried out by two individuals and evaluated by a third party for accuracy. (2) We use the augmented and altered captions to generate reference images with a suite of Text-2-Image models, including DALL-E 2 and Stability Diffusion XL for 4 variants of styles. (3) We implement an additional quality check process for all generated images to eliminate and regenerate images that might contain unsafe or counterintuitive content. We employ BLIP2 [30] to filter out generated images that have lower semantic similarity with augmented captions than 0.2 and conduct a manual sanity check to control the image quality.
A.1.3 Statistics
@@
Reference Image Style | #Queries | #Videos | #With Refinement Texts | #Without Refinement Texts |
scribble | 1546 | 1515 | / | 5 |
cinematic | 1532 | 1502 | 1445 | 5 |
cartoon | 1532 | 1501 | 1444 | 5 |
@@
[…]
@@
A.1.4 Details of Deleted Data
“A graph depicts penis size.” (qid: 9737)
“People mess with the bull statues testicles.” (qid: 7787)
“People butcher meat from a carcass.” (qid: 4023)
A.2 Benchmark Details
A.2.1 Model Comparison
Tab. 6 compares our selected baseline models. Query encoder denotes the text encoder of each model to encode natural language queries. Source represents the modalities of the source data, while V and A refer to “Video” and “Audio” respectively. All models have been fine-tuned on QVHilight.
@@ A.2.1 Model Comparison
SeViLA (2023) | ViT-G | CLIP Text | Multimodal LLM (BLIP2) | V |
A.2.2 Prompt Engineering
Since the performance may highly depend on the wording in a prompt, we use 3 different prompts for Cap and Sum adaptation methods. In Tab. 7, the prompts are divided into “Prompts For Style cartoon/cinematic/realistic” and “Prompts for scribble”. This distinction arises because refining scribble images with complementary texts involves adding new details, slightly differing from other scenarios. Despite this minor variation, the prompt style remains consistent, simulating 3 different user query styles.

### v3

@@ 4.1 Multimodal Query Adaptation
VQ-Enc
Moment-DETR (2021) 12.55 5.69 13.38 6.59 14.36 6.01 14.88 6.53
QD-DETR (2023) 15.91 9.12 14.88 8.62 13.90 8.49 14.62 8.36
QD-DETR (2023) 15.65 10.03 12.60 6.79 12.34 6.72 12.34 7.44
EaTR (2023) 19.86 13.00 19.91 12.99 21.15 13.45 21.48 13.38
CG-DETR (2023) 22.90 13.00 24.93 13.58 23.24 13.12 24.74 14.23
TR-DETR (2024) 17.92 11.19 17.36 11.10 15.14 9.86 15.60 9.53
UMT (2022) 5.43 2.85 4.77 2.09 5.22 2.35 4.57 2.42
UniVTG (2023) 21.93 13.00 23.89 13.64 22.78 13.19 22.52 12.79
UVCOM (2023) 17.08 9.77 16.78 10.97 17.36 11.68 17.10 11.23
MQ-Cap
@@
TimeChat (2024) 6.58 (± 0.1) 2.76 (± 0.5) 7.38 (± 1.1) 3.39 (± 0.8) 7.51 (± 0.9) 3.63 (± 0.8) 5.73 (± 1.2) 4.49 (± 3.3)
VTimeLLM (2024) 16.95 (± 1.4) 7.40 (± 0.1) 19.19 (± 0.8) 7.8 (± 0.3) 20.23 (± 0.4) 8.29 (± 0.3) 20.53 (± 1.5) 8.11 (± 0.5)
MQ-Sum
+ SUIT
Moment-DETR (2021) 48.59 (± 0.9) 31.85 (± 0.7) 48.27 (± 0.6) 31.31 (± 0.4) 47.58 (± 0.5) 31.52 (± 0.5) 47.25 (± 0.2) 30.83 (± 0.6)
QD-DETR (2023) 55.27 (± 0.5) 39.86 (± 0.4) 53.45 (± 0.6) 37.94 (± 0.3) 53.36 (± 0.3) 38.39 (± 0.6) 53.79 (± 0.5) 38.92 (± 0.1)
QD-DETR†(2023) 55.20 (± 0.5) 39.82 (± 0.7) 54.60 (± 0.4) 40.44 (± 0.6) 54.28 (± 0.4) 40.31 (± 0.6) 53.52 (± 0.8)
38.97 (± 0.1)
EaTR (2023) 53.63 (± 0.8) 39.23 (± 0.5) 50.63 (± 0.4) 37.40 (± 0.6) 51.67 (± 0.5) 38.50 (± 0.4) 50.78 (± 0.4) 37.19 (± 0.5)
CG-DETR (2023) 55.83 (± 0.6) 41.41 (± 0.3) 55.42 (± 0.8) 39.88 (± 0.6) 56.37 (± 0.8) 41.14 (± 0.6) 55.47 (± 0.9) 40.17 (± 0.5)
TR-DETR (2024) 58.85 (± 0.4) 43.08 (± 0.4) 57.19 (± 0.2) 41.31 (± 0.4) 57.35 (± 0.5) 41.92 (± 0.9) 57.39 (± 0.4) 42.64 (± 0.3)
UMT†(2022) 49.71 (± 0.3) 35.10 (± 0.3) 50.01 (± 0.8) 35.16 (± 0.6) 50.25 (± 0.6) 35.18 (± 0.5) 49.85 (± 0.4) 34.60 (± 0.7)
UniVTG (2023) 51.26 (± 0.4) 34.07 (± 0.7) 49.36 (± 0.3) 33.24 (± 0.5) 51.0 (± 0.5) 34.4 (± 0.7) 50.65 (± 0.6) 33.48 (± 0.6)
UVCOM (2023) 55.33 (± 0.4) 42.03 (± 0.7) 55.48 (± 0.2) 41.66 (± 0.1) 55.43 (± 0.4) 41.88 (± 0.4) 54.43 (± 0.4) 41.30 (± 0.3)
4.2 SUIT: Surrogate Fine-tuning on Pseudo-MQs
Fine-tuning MLLMs on the task of summarizing MQs could counteract the impact of different prompt selections and mitigate the distribution shift between original NLQs and generated NLQs. However, an underlying challenge for fine-tuning lies in the lack of training data for MQ-based localization. Compared to establishing an evaluation testbed, the larger-scale training data is more time and labor-intensive. Besides, synthetic training data could pose risks of overfitting on generation bias and artifacts in the model, which is supposed to be avoided.
To overcome this challenge, we propose a novel strategy, SUrrogate FIne-Tuning (SUIT) on pseudo-MQs, to alleviate the training data issue.
As illustrated in Fig. 4, SUIT consists of 2 steps:
Pseudo-MQ Generation Pipeline To deal with the insufficient training data problem, we propose leveraging the abundant image-text datasets like Flickr30K [39] and COCO [45] to generate pseudo-MQs. We automate this generation process by leveraging GPT3.5 to convert each caption in the datasets to a pair of a “forged” caption and a refinement text that reflects the forge. As a result, the original image and the refinement text constitute a pseudo-MQ that is equivalent to a forged caption semantically.
Surrogate Fine-tuning on Psuedo-MQs We further utilize generated pseudo-MQs as inputs and instruct MLLMs to generate a summarization as MQ-Sum. Distorted captions are used as supervision to fine-tune the model with the next-token prediction loss and the PEFT approach as a surrogate training task. Then, we can transfer the fine-tuned MLLMs to our ICQ-Highlight dataset for evaluation.
4.3 Backbone Model Selection
We have selected and benchmarked 12 models specifically designed for video event localization with NLQs. Particularly, we categorize the selected models as follows and compare the models in different dimensions in the Appendix: (1) Specialized models use natural language as a semantic query and are targeted at video moment retrieval tasks. We have selected a series of these models including Moment-DETR[42], QD-DETR[60], EaTR[35], CG-DETR[59], and TR-DETR[71]; (2) Unified frameworks are aimed to solve multiple video localization tasks within one model, such as moment retrieval, highlight detection, and video summarization. We have selected UMT[54], UniVTG[44], and UVCOM[83] as strong baselines; (3) LLM-based Models features the power of Large Language Models, which prove to be a powerful and general head for varied video tasks. We have selected SeViLA [91], TimeChat [66], and VTimeLLM [33] as representatives of LLM-based models. We apply different MQA methods on top of the pre-trained model checkpoints that have been fine-tuned on the original QVHighlights dataset.
5 Experiments and Analysis
In this section, we attempt to answer the following questions: (1) Can and how well MQs effectively localize events in videos? (2) Can varied styles of reference images and refinement texts impact the results?
5.1 Experimental Setup
Implementation We employ LLaVA-mistral-1.6 [50, 51] as a strong MLLM in MQ-Cap, MQ-Sum (with and without SUIT) and GPT-3.5 as a reviser in our MQ-Cap adaptation. We believe that the performance of these models is representative of the SOTA capabilities of MLLMs and is fairly compared across different MQA methods. For VQ-Enc, we utilize the corresponding CLIP Visual Encoder, as all models typically employ the CLIP Text Encoder for text query encoding. In this adaptation method, we omit refinement texts and only use the reference image. In MQ-Sum(+SUIT), we construct our pseudo-MQs with training data from Flickr30K and COCO and implement LoRA [32] as a common PEFT method with rank , alpha , and a learning rate of on language model of LlaVA. More implementation details about datasets and training can be found in the Appx. B.1.
Evaluation Metrics We evaluate models on our new testbed ICQ-Highlight. For evaluation, we report both Recall R@1 with IoU thresholds and , mean Average Precision with IoU threshold 0.5 and the average over multiple IoU thresholds [0.5:0.05:0.95] as standard metrics for video moment retrieval and localization [42, 91], where IoU (Intersection over Union) thresholds determine if a predicted temporal window is positive.
@@ 5.1 Experimental Setup
5.2 Results & Analysis
We present the pairwise performance of 12 models combined with 4 adaptation methods on ICQ in Tab. 4.1 and Tab. C.1 in Appx. C.1. For MQ-Cap and MQ-Sum methods, we have conducted multiple runs with different prompts and reported the average performance and standard deviation.
How do Video Event Localization with MQs work on different image styles? Firstly, we aim to draw a key conclusion from the results. We find all adaptation methods perform consistently across different styles and therefore suggest that they could understand the MQs well, particularly for styles including cartoon, cinematic, and realistic; the model performance is close to each other. For scribble, all models show marginally worse performance, and even both MQ-Cap and MQ-Sum methods have a more significant standard deviation, which reflects that it is heavily influenced by the prompts. This can be explained by the fact that scribble images are more minimal and abstract in semantics and more challenging to interpret. Surprisingly, in spite of being more abstract and simpler, the model performance on scribble reference images is close to other reference image styles. This demonstrates the potential of using scribble as MQs in real-world video event localization applications like video search.
Which is the best MQA method? Among all the training-free methods, we find that MQ-Cap can achieve the best performance and is more robust to different prompts compared to other adaptation methods by an average margin of on all styles. We observe that both utilizing MLLMs for captioning reference images, MQ-Sum suffers more than MQ-Cap adaptation regarding performance and is more sensitive to prompts for all reference styles, which can be observed from the higher standard deviation, showing asking MLLMs to caption and summarize the refinement texts is less controllable. To conclude, captioning images is still a golden method since MLLMs and LLMs are powerful enough to generate faithful captions.
Notably, MQ-Sum(+SUIT) shows a non-marginal improvement () and more stable performance across all backbone models. This proves the efficacy and transferability of our SUIT strategy. To verify our motivation that training-free MQA can output uncontrollable text queries that have a distribution shift from the original NLQs on which the backbones are trained, we visualize the embeddings of original NLQs and adapted MQs in Fig. 6 with t-SNE [76]. It shows that original NLQs have similar distributions as MQ-Sum(+SUIT) other than the other 2 training-free methods for all different image styles.
However, the performance gap between our MQ setting and the original NLQ benchmark (refer to Appx. C.5) is still remarkable, which shows that the query semantics are more or less distorted across modalities.
Across different backbone models, we find that models that perform well in one adaptation method tend to perform well in others. For example, UVCOM and TR-DETR consistently show high performance across MQ-Cap, MQ-Sum, and VQ-Enc methods. We observe that more recent models keep their outperforming performance on our ICQ. Latest models, including UVCOM, TR-DETR, and CG-DETR, tend to perform better across different adaptation methods and reference image styles. In contrast, older models like Moment-DETR consistently show lower performance. LLM-based models cannot compete with other specialized models without exception; this aligns with their subpar performance on NLQ-based benchmarks [92, 33, 66]. In the next section, we find that model performance on ICQ highly correlates with that on natural language query-based benchmark QVHighlights. This shows that (1) our multimodal queries share semantics with the original benchmark; (2) the adaptation methods and models could understand semantics from multimodal queries.
5.3 Ablation Studies
Besides the benchmark, we conduct additional studies for other intriguing questions in this section and in Appx. C.1.
Do Artifacts in synthetic reference images distort the conclusion? The artifacts in our generated data are inevitable even with the best commercial Text-to-Image models so far. To understand the impact of generated images’ artifacts on model evaluation, we conduct a controlled experiment by collecting a subset of MQs by crawling similar images via the Google image search engine. Each image in this retrieved subset has a corresponding generated reference image in a subset of ICQ-Highlight. The retrieval criterion is that retrieved images should be as similar as possible to the generated images in semantics/style/details so that the generation artifacts are the only control variable. The final subset comprises 84 samples from 4 styles. We compare the model performance on and . Our pre-assumption is that if generation artifacts degrade the model performance largely, then should perform better than . Otherwise, should perform close to . As shown in Fig. 5, model performance on is close to in general. This shows that generation artifacts do not skew our findings largely, and our benchmark is still generalizable.
Importance of Refinement Texts To assess the impact of refinement texts on video event localization using MQs, we have evaluated model performance using only reference images as queries, omitting refinement texts. We employ the MQ-Cap adaptation without a modifier for integrating refinement texts. As shown in Tab. 2, we present the model performance and their relative performance drop in percentage compared to those with refinement texts. Models have different scales of performance drop, which indicates that refinement texts help refine the semantics of reference images and localize the events. Additionally, we observe that for scribble images, the performance drop is less pronounced compared to other styles in that these images are inherently minimalistic and less reliant on details.
@@ 6 Conclusion
Societal Impacts Using multimodal semantic queries for video event localization brings prospects in real-world applications, such as assisting illiterate, pre-literate, or non-speakers in cross-lingual situations, as it allows them to interact with videos through images as a more accessible and convenient approach.
In this work, we introduce a new benchmark, ICQ, marking an initial step towards using multimodal semantic queries for video event localization. We have found that our proposed MQA and SUIT methods can accommodate conventional models to MQs effectively, serving as effective baselines for this novel setting. Our findings confirm that using MQs for video event localization is practical and feasible. Nonetheless, the field remains open to innovative model architectures and training paradigms for MQs. We believe our work paves the way for real-world applications that leverage MQs to interact with video content.
Appendix
@@
Additional information about the dataset ICQ-Highlight and licenses for the datasets and models we have used;
@@
Appendix A Dataset: ICQ-Highlight
A.1 License
@@
A.2 Construction Pipeline
We base our model on the original annotation from QVHighlights [42]. The whole pipeline, as shown in Fig. 7 consists of (1) annotation: We further conduct a quality check on the annotations in the original dataset and filter out a few samples (details can be found in Sec. A.4). In order to generate more relevant reference images, we manually augment the original captions by adding new visual details based on three frames extracted from the raw videos. To introduce refinement texts, we purposely alter certain details of the captions to generate a new one. All annotations are carried out by two individuals and evaluated by a third party for accuracy. (2) We use the augmented and altered captions to generate reference images with a suite of Text-2-Image models, including DALL-E 2 and Stability Diffusion XL for 4 variants of styles. (3) We implement an additional quality check process for all generated images to eliminate and regenerate images that might contain unsafe or counterintuitive content. We employ BLIP2 [43] to filter out generated images with lower semantic similarity with augmented captions than 0.2 and conduct a manual sanity check to control the image quality.
Data Curation and Quality check Image generation can suffer from significant imperfections in terms of semantic consistency and content safety. To address these issues, we implement a quality check in 2 stages: (1) We calculate the semantic similarity between the generated images and the text queries using BLIP2 [43] encoders, eliminating samples that score lower than 0.2; (2) We perform human sanity check to replace images that are: i) semantically misaligned with the text, ii) mismatched with the required reference image style, iii) containing sensitive or unpleasant content (e.g., violent, racial, sexual content), counterintuitive elements, or noticeable generation artifacts.
A.3 Statistics
The dataset comprises 1515 videos and 1546 test samples on average for each style. The exact numbers may vary slightly across styles and are provided in the Appendix.
@@ A.3 Statistics
#Queries | #Videos |
| scribble | 1546 | 1515 | / | 5 | ||||||
| cinematic | 1532 | 1502 | 1445 | 5 | ||||||
| cartoon | 1532 | 1501 | 1444 | 5 | ||||||
@@
| Object | Action | Relation | Attribute | Environment | Others | |
@@
| SeViLA (2023) | ViT-G/14 | CLIP Text | Glan-T5 XL (3B) [17] | V |
| TimeChat (2024) | ViT-G/14 + Video Q-Former | LLaMA tokenizer | LLaMA-2 7B [26] | V |
| VTimeLLM (2024) | ViT-L/14 | LLaMA tokenizer | Vicuna v1.5 (7B) [16] | V |
A.4 Details of Deleted Data
“A graph depicts penis size.” (qid: 9737)
“People mess with the bull statues testicles.” (qid: 7787)
“People butcher meat from a carcass.” (qid: 4023)
Appendix B Benchmark Details
In this section, we list the details of our selected backbone models, the implementation of our training-free MQA methods, and SUIT strategy.
B.1 Implementation Details
Automatic Pseudo-MQs Construction
We build the pseudo-MQ dataset from image-text datasets Flickr30K and COCO. We generate captions for the COCO dataset with BLIP-2 [43]. To forge the original captions, we employ GPT3.5 to process the pure-text captions of each image with the prompts shown in Tab. 10. For each sample, we randomly select one template and refinement text type to generate a forged caption and the corresponding forged part as a refinement text. In total, we construct a pseudo-MQ dataset with samples for training and samples for validation.
Implementation of SUIT
We apply LoRA to all linear layers in the language model of LLaVA-mistral-1.6 with and with one epoch on the full dataset. The training takes up to 16 hours on a single NVIDIA A40 GPU.
B.2 Model Comparison
Tab. 5 compares our selected baseline models. The query encoder denotes the text encoder of each model used to encode natural language queries. Source represents the modalities of the source data, while V and A refer to “Video” and “Audio” respectively. All models have been fine-tuned on QVHilights.
B.3 Prompt Engineering
Since the performance may highly depend on the wording in a prompt, we use 3 different prompts for MQ-Cap and MQ-Sum adaptation methods. In Tab. 6, the prompts are divided into “Prompts For Style cartoon/cinematic/realistic” and “Prompts for scribble”. This distinction arises because refining scribble images with complementary texts involves adding new details, slightly differing from other scenarios. Despite this minor variation, the prompt style remains consistent, simulating 3 different user query styles.
For MQ-Sum(+SUIT), we use the same prompts as MQ-Sum in the parameter-efficient fine-tuning with LoRA.